What is a+(a-2)=(a+3)-(a+1)
Found 2 solutions by josgarithmetic, addingup:
Answer by josgarithmetic(39617)
Answer by addingup(3677)

a+(a-2)=(a+3)-(a+1)
= a+a-2 = a+3-(a+1)
= 2a-2 = a+3+((-1)a-1)Note: nothing changed here, the -1 stands for -a. When you multiply the -1 using the distributive properties of multiplication, your get:
-1*a and -1*-1 = -1a and +1, or simply -a+1. Let's continue. Remove unnecessary parenthesis:
= 2a-2 = a+3+(-1)a-1 Add 3 and -1
= 2a-2 = a+2+(-1)a Now a-1a = 0, let's rewrite:
= 2a-2 = 2
= 2a-2-(-2) = 2-(-2)
= 2a = 2+2
= 2a = 4
= a = 4/2
= a = 2
----------------------
Check, see if I got an equality:
a+(a-2)=(a+3)-(a+1)
2+(2-2) = (2+3)-(2+1)
2 = 5-3; 2 = 2 Correct